\begin{thebibliography}{10}






\bibitem{Avramov}
 L. L. Avramov, \emph{Modules of finite virtual projective dimension,} Invent. Math. \textbf{96} (1989), no. 1, 71–101. MR 981738

 \bibitem{Peeva}
Luchezar L. Avramov, Vesselin N. Gasharov, and Irena V. Peeva, \emph{Complete intersection dimension,} Inst. Hautes \'{E}tudes Sci. Publ. Math. (1997), no. 86, 67–114. MR 1608565

\bibitem{SBrodman}
M. P. Brodmann and R. Y. Sharp, \emph{Local cohomology: an algebraic introduction with geometric applications,} Cambridge Studies in Advanced Mathematics, vol. 60,
Cambridge University Press, Cambridge, 1998. MR 1613627

 \bibitem{Bruns}
Winfried Bruns and J\"urgen Herzog, \emph{Cohen-Macaulay rings,} Cambridge Studies in Advanced Mathematics, vol. 39, Cambridge University Press, Cambridge, 1993. MR 1251956

\bibitem{Eisenbud}
David Eisenbud, \emph{Homological algebra on a complete intersection, with an application to group representations,} Trans. Amer. Math. Soc. \textbf{260} (1980), no. 1, 35–64. MR 570778

\bibitem{Gulliksen}
Tor H. Gulliksen, \emph{A change of ring theorem with applications to Poincar\'e series and intersection multiplicity,} Math. Scand. \textbf{34} (1974), 167–183. MR 364232

\bibitem{HCCMM}
 Tony J. Puthenpurakal, \emph{Hilbert-coefficients of a Cohen-Macaulay module,} J. Algebra \textbf{264} (2003), no. 1, 82–97. MR 1980687

 \bibitem{HCMCMPart1}
\bysame, \emph{The Hilbert function of a maximal Cohen-Macaulay module,} Math. Z. \textbf{251} (2005), no. 3, 551–573. MR 2190344

\bibitem{Quasipolynomial}
\bysame, \emph{Growth of Hilbert coefficients of syzygy modules,} J. Algebra \textbf{482} (2017), 131–158. MR 3646287

\bibitem{Part1}
\bysame, \emph{Ratliff-Rush filtration, regularity and depth of higher associated graded modules. I,} J. Pure Appl. Algebra \textbf{208} (2007), no. 1, 159–176. MR 2269837

\bibitem{cxM}
\bysame, \emph{On the Stable category of maximal Cohen-Macaulay modules over Gorenstein rings,} arXiv preprint arXiv:2107.05237 (2021).

\bibitem{TONYT}
\bysame, \emph{sub-functor for Ext and Cohen-Macaulay associated graded modules with bounded multiplicity,} Trans. Amer. Math. Soc. \textbf{373} (2020), no. 4, 2567–2589. MR 4069228

\bibitem{JDSally}
Judith D. Sally, \emph{Numbers of generators of ideals in local rings,} Marcel Dekker, Inc., New York-Basel, 1978. MR 485852

\bibitem{NVT3}
Ng\^o Viet Trung, \emph{Reduction exponent and degree bound for the defining equations of graded rings,} Proc. Amer. Math. Soc. \textbf{101} (1987), no. 2, 229–236. MR 902533


\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{amsmath,amsfonts, mathtools,amssymb,latexsym,amssymb,amsthm,adjustbox,mathrsfs,amsbsy, bm,tikz-cd,indentfirst, makecell,graphicx, verbatim, calc, enumerate,xy}
\usepackage[top=1in,bottom=1.3in,left=1.1in,right=1.1in]{geometry}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{remark}[theorem]{Remark}

\newtheorem{question}[theorem]{Question}

\newtheorem{point}[theorem]{}

\newcommand{\m}{\mathfrak{m}}

\newcommand{\R}{\mathcal{R}}
\newcommand{\F}{\mathcal{F}}
\newcommand{\Z}{\mathbb{Z}}
\newcommand{\N}{\mathbb{N}}
\newcommand{\wt}{\widetilde}
\newcommand{\n}{\mathfrak{n}}

\newcommand{\cx}{\operatorname{cx}}

\newcommand{\Ext}{\operatorname{Ext}}
\newcommand{\Tor}{\operatorname{Tor}}

\usepackage{graphicx}
\usepackage{subfigure}
\usepackage{epsfig}
\usepackage{epstopdf}
\usepackage{float}
\theoremstyle{plain}

\usepackage{xcolor}
\usepackage{titlesec}

\titleformat{name=\section}{}{\thetitle.}{0.5em}{\centering\scshape\large}

\begin{document}
\title{\large \textbf{Asymptotic behavior of invariants of syzygies of maximal Cohen-Macaulay modules }}
\author{\textsc{Tony J. Puthenpurakal}}
\email{tputhen@gmail.com}
\address{Department of Mathematics, IIT Bombay, Powai, Mumbai 400 076, India}
\author{\textsc{Samarendra Sahoo}}
\email{204093008@iitb.ac.in}

\address{Department of Mathematics, IIT Bombay, Powai, Mumbai 400 076, India}
\date{\today}

\subjclass{Primary 13A30, 13C14, 13D40, Secondary  13D02, 13D07, 13D45}
\keywords{Complete intersection rings, Complexity of modules, Associated graded rings, Castelnuovo-Mumford regularity, Ulrich modules, superficial element, Hilbert coefficients}

\begin{abstract} Let $(A,\m)$ be a complete intersection ring of codimension $c\geq 2$ and dimension $d\geq 1$. Let $M$ be a finitely generated maximal Cohen-Macaulay $A$-module. Set $M_i=\text{Syz}^A_{i}(M)$. Let $e^{\m}_i(M)$ be the $i$-th Hilbert coefficient of $M$ with respect to $\m$. We prove for all $i\gg0$, the function $i\mapsto e^{\m}_j(M_i)$ is a quasi-polynomial type with period $2$ and degree $\text{cx}(M)-1$ for $j=0,1$, where $\text{cx}(M)$ is the complexity of $M.$ For $\text{cx}(M)=2,$ we prove  $$\lim_{n\to \infty}\dfrac{e^{\m}_1(M_{2n+j})}{n}\geq \lim_{n\to \infty}\dfrac{e^{\m}_0(M_{2n+j})}{n}-\lim_{n\to \infty}\dfrac{\mu(M_{2n+j})}{n}$$ for $j=0,1$. When equality holds, we prove that the Castelnuovo-Mumford regularity of the associated graded ring of $M_i$ with respect to the maximal ideal $\mathfrak{m}$ is bounded for all $i\geq 0$.

\end{abstract}

\maketitle

\section{Introduction}
Let $(A,\mathfrak{m})$ be a Noetherian local ring with residue field $k=A/\mathfrak{m}$. Let $I$ be a $\mathfrak{m}$-primary ideal and $M$ be a finitely generated $A$-module. We denote the associated graded ring of $A$ with respect to the ideal $I$ as $G_I(A)=\bigoplus_{n\geq 0}I^n/I^{n+1}$, and the associated graded module of $M$ with respect to $I$ as $G_I(M)=\bigoplus_{n\geq 0}I^nM/I^{n+1}M$, considered as a graded $G_I(A)$-module. Let $\ell(E)$ denotes the length of an $A$-module $E$.

Let $M$ be an $A$-module of dimension $r$. The Hilbert series of $M$ is given by the formula $$H_{M,I}(z)=\sum_{i\geq 0}\ell(I^iM/I^{i+1}M)z^i=\frac{h^I_M(z)}{(1-z)^r},$$ where $h^I_M(z)\in \Z[z]$ and is called the $h$-polynomial of $M$. The integer $e^I_i(M)=h^{(i)}_{M}(1)/i!$ is referred to as the $i$-th Hilbert coefficient of $M$ with respect to $I$, where $h^{(i)}_M(z)$ is the $i$-th derivative of $h_M(z)$. When $I=\mathfrak{m}$, we write $G(A)=G_I(A)$, $G(M)=G_I(M)$ and $e_i(M)=e^I_i(M)$.

There has been extensive research on the behavior of the Betti numbers of the Cohen-Macaulay module over complete intersection rings. For instance, see \cite{Avramov}, \cite{Peeva}, and \cite{Eisenbud}. Set $\beta_i(M)$ as the $i$-th Betti number of $M$, and $\mu(M)$ as the cardinality of the minimal set of generators of $M$. Note that $\beta_i(M)=\mu(M_i)$ for all $i\geq 0$. One interesting result due to Avramov (\cite{Avramov}, Theorem 4.1) is that the function $i\mapsto \beta_i(M)$ is a quasi-polynomial (for the definition of quasi-polynomial, see \ref{qasi}) type with period 2 for sufficiently large $i$. Furthermore, the polynomials $P_0$ and $P_1$ have the same degree and leading coefficient. Similar questions can be posed regarding the behavior of other invariants of $M_i$ for sufficiently large $i.$

Let $M_i=\text{Syz}^A_i(M)$ for all $i\geq 1$, representing the $i$-th syzygy of $M$ in its minimal free resolution, and for convenience, set $M_0=M$. In \cite{Quasipolynomial}, the first author focused on the Hilbert coefficients of the syzygies of the modules and established that for a maximal Cohen-Macaulay module $M$ over complete intersection rings (see \cite{Quasipolynomial}, 2.4), the function $i\mapsto e^I_i(M_i)$ is a quasi-polynomial type with period 2 and degree at most $\text{cx}(M)-1$ for all $0\leq i\leq 2$. In this paper, we explore the degree of the function $i\mapsto e_t(M_i)$ for $t=0,1$ and reg $G(M_i)$ (for the definition of regularity, see \ref{regL}) for sufficiently large values of $i$ over complete intersection rings.

In the first part of this paper (Section $3$), we will address the following questions:

\begin{question}
\label{que}
Let $(A,\mathfrak{m})$ be a complete intersection ring and $M$ be a maximal Cohen-Macaulay $A$-module.
\begin{enumerate}
    \item[{(a)}] When does the function $i\mapsto e_t(M_i)$ attain its maximal degree for $t= 0,1$?
    \item[{(b)}] When do the polynomials $P_0$ and $P_1$ have the same degree?
\end{enumerate}
\end{question}
For the leading coefficients of $P_0$ and $P_1$, we have both positive and negative answers for the function $i\mapsto e_t(M_i)$ for $t=0,1$. For example,
\begin{enumerate}
    \item Let $(A,\m)=(k[|x,y|]/(x^2y^2),(\overline{x},\overline{y}))$, where $k$ is a field and $M=A/(\overline{x^2})$. Note that $M$ is a maximal Cohen-Macaulay $A$-module. It is easy to check that $e_0(M_i)=2$ and $e_1(M_i)=1$ for all $i\geq 0.$
    \item Let $M$ be a maximal Cohen-Macaulay module with $e_0(M)=\mu(M)$ (in particular, $M$ is Ulrich) over a hypersurface ring with $e_0(A)>2$. Then, $e_0(M_{2i})\neq e_0(M_{2i+1})$ for all $i\geq 0$ (see \cite{HCMCMPart1}, Theorem 2, Remark 4.10).
    \item Let $(A,\m)=(k[|x,y|]/(xy^2),(\overline{x},\overline{y}))$ and $M=A/(\overline{x})$. Note that $M$ is a maximal Cohen-Macaulay $A$-module. It is easy to check that $e_1(M_{2i})=0$ and $e_1(M_{2i+1})=1$ for all $i\geq 0.$
\end{enumerate}
However, we are unable to provide sufficient conditions for the leading coefficients to be equal. Additionally, the Question \ref{que} remains open for $t\geq 2$.

First, we will present the following result, which provides a positive response to Question \ref{que} for $t=0$ and for an arbitrary $\m$-primary ideal, which is not very difficult to prove.
\begin{proposition}
\label{1st}
    Let $(A,\m)$ be a complete intersection ring and $I$ a $\m$-primary ideal. Let $M$ be a maximal Cohen Macaulay module. Then the function $i\mapsto e^I_0(M_i)$ is a quasi-polynomial type with period $2$ and degree $\operatorname{cx}(M)-1$ for all $i\gg0.$
\end{proposition}

In order to prove Question \ref{que} for the first Hilbert coefficient, additional conditions are required for both rings and the ideal. For example, let $I=(x_1,\ldots ,x_d)$ be an ideal of $A$ generated by a maximal $A$- regular sequence. Then $x_1,\ldots ,x_d$ is a maximal $M_i$-regular sequence for all $i\geq 0.$ Therefore, for all $i\geq 0$, the Hilbert series of $M_i$ is $H_{M_i,I}(z)=\dfrac{h^I_{M_i}(z)}{(1-z)^d},$ where $h^I_{M_i}(z)=\ell(M_i/IM_i)$ (see \cite{Bruns}, Theorem 1.1.8). Hence $e^I_1(M_i)=0$ for all $i\geq 0.$ We prove the following for the maximal ideal.

\begin{theorem}
\label{2nd}
     Let $(A,\m)$ be a complete intersection ring of codimension $c\geq 2$ and dimension $d\geq 1$. Let $M$ be a maximal Cohen Macaulay $A$-module. Set $M_i=\operatorname{Syz}^A_i(M)$ for all $i\geq 0.$ Then the function $i\mapsto e_1(M_i)$ is a quasi-polynomial type with period $2$ and degree $\operatorname{cx}(M)-1$ for all $i\gg0.$
\end{theorem}

Next, we will compare the leading coefficients of the functions $i \mapsto \mu(M_i)$, $i \mapsto e_0(M_i)$, and $i \mapsto e_1(M_i)$ for sufficiently large $i$.

 \begin{lemma}
 \label{3rd}
     Let $(A,\m)$ be a complete intersection ring of codimension $c\geq 2$ and dimension $d\geq 1$. Let $M$ be a maximal Cohen Macaulay $A$-module. Then for $j=0,1$ $$\lim_{n\to \infty}\dfrac{e_1(M_{2n+j})}{n^{\operatorname{cx}(M)-1}}\geq \lim_{n\to \infty}\dfrac{e_0(M_{2n+j})}{n^{\operatorname{cx}(M)-1}}-\lim_{n\to \infty}\dfrac{\mu(M_{2n+j})}{n^{\operatorname{cx}(M)-1}}.$$
\end{lemma}

Let $M$ be an $A$-module with dimension $d$. One interesting question is how the invariants $a_j(M_i)$ (see \ref{regL}) change for large values of $i$ and for $0\leq j\leq d$. In (\cite{Quasipolynomial}, Theorem 8.1), the first author established that the set $\{a_{d-1}(G(M_n))/n^{\operatorname{cx}(M)-1}\, |\, n\geq 1\}$ is bounded. However, in this paper, we prove an even stronger result, which is as follows:

 \begin{theorem}
 \label{5th}
     Let $(A,\m)$ be a complete intersection ring of codimension $c\geq 2$ and dimension $d\geq 1$. Let $M$ be a maximal Cohen Macaulay $A$-module. If $\operatorname{cx}(M)=2$ and $$\lim_{n\to \infty}\dfrac{e_1(M_{2n+j})}{n}= \lim_{n\to \infty}\dfrac{e_0(M_{2n+j})}{n}-\lim_{n\to \infty}\dfrac{\mu(M_{2n+j})}{n}$$ for $j=0,1$. Then $\text{reg }G(M_n)$ is bounded for all $n\gg0.$
\end{theorem}

We will now provide a brief overview of this paper. In Section 2, we will present some necessary preliminary results. Section 3 will outline the proofs of Theorem \ref{1st} and Theorem \ref{2nd}. In Section 4, we will present evidence for Lemma \ref{3rd}. Finally, in Section 5, we will demonstrate Theorem \ref{5th} using Lemma \ref{3rd}.

\section{Preliminaries}
Throughout this paper all rings considered are Noetherian and all modules considered, unless stated otherwise, are finitely generated.

\begin{point}
Base change \\
\normalfont
   Let $\phi:(A,\mathfrak{m})\to (A',\mathfrak{m}')$ be a local ring homomorphism. Assume $A'$ is a faithfully flat $A$-algebra with $\mathfrak{m}A=\mathfrak{m}'$. Let $M$ be an $A$-module. Set $M'=M\otimes A'$. In these cases, it can be shown that
\begin{enumerate}
    \item $\ell_A(M)=\ell_{A'}(M')$.
    \item $H(M,n)=H(M',n)$ for all $n\geq 0.$
    \item dim $M=$ dim $M$ and $\text{ depth}_A M= \text{ depth}_{A'} M'$.
    \item depth $G(M)=\text{ depth } G(M')$.
    \item $A$ is a local complete intersection iff $A'$ is a local complete intersection.
\end{enumerate}
\end{point}
The specific base changes we do are the following:
\begin{enumerate}
    \item[(i)] $A'=\hat{A}$ the completion of $A$ with respect to the maximal ideal.
    \item[(ii)] $A'=A[[X]]_S$, where $S=A[[X]]\setminus \m A[[X]].$ The maximal ideal of $A'$ is $\m '=\m A'$ and the residue field of $A'$ is $K=k((X)).$
\end{enumerate}
Thus, we can assume that the ring A is complete and the residue field is uncountable.

\begin{point}
The Rees module $L(M)$\\
\label{hpolyn}
      \normalfont
     Let $(A,\m)$ be a local ring and $M$ be an $A$-module. Define  $L(M)=\bigoplus_{n\geq 0}M/\m^{n+1}M$. Let $\R=A[\m t]$ be Rees ring and
  $\mathcal{R}(M)=\bigoplus_{n\geq 0}\m^nMt^n$ be Rees module of $M$. The Rees ring $\mathcal{R}$ is a subring of $A[t]$. So $A[t]$ is an $\mathcal{R}$ module. Therefore $M[t] = M\otimes_A A[t]$ is an $\R$-module. The exact sequence $$0\to \mathcal{R}(M)\to M[t]\to L(M)(-1)\to 0 $$ defines an $\mathcal{R}$ module structure on  $L(M)(-1)$ and hence on $L(M)$ (for more details see (\cite{Part1}, definition 4.2)). Note that $L(M)$ is not a finitely generated $\R$-module.
\end{point}

\begin{point}
    \normalfont
The Hilbert-Samuel function of the module $M$ with respect to the ideal $\mathfrak{m}$ is defined as the function $n \mapsto \ell(M/\mathfrak{m}^{n+1}M)$. It is well-known that for large values of $n$, this function can be expressed as a polynomial $P_M(n)$ of degree equal to the dimension of $M$ (say $r$), known as the Hilbert-Samuel polynomial of $M$. The polynomial can be written in the form $$P_M(n)=e_0(M)\binom{n+r}{r}-e_1(M)\binom{n+r-1}{r-1}+\ldots +(-1)^re_r,$$ where $e_0(M), \ldots , e_r(M)$ are the Hilbert coefficients of $M$.

 It can be easily seen that the formal power series $\sum_{n\geq 0}\ell{(M/\m^{n+1}M)}z^n$ represents a rational function $$\sum_{n\geq 0}\ell{(M/\m^{n+1}M)}z^n=\dfrac{h_M(z)}{(1-z)^{r+1}},$$ i,e $H(L(M),z)=\dfrac{h_M(z)}{(1-z)^{r+1}}.$  \end{point}

\begin{point}
\label{L_1M}
    \normalfont
    Set $L_1(M)=\bigoplus_{n\geq 0}\Tor^{A}_1(M,A/\m^{n+1}).$ Note that $L(A)(-1)\otimes M=L(M)(-1).$ By tensoring the exact sequence $0\to \mathcal{R}\to A[t]\to L(A)(-1)\to 0$ with $M$, we get $$0\to L_1(M)(-1)\to \mathcal{R}\otimes M\to M[t]\to L(M)(-1)\to 0.$$ Thus $ L_1(M)(-1)$ is a $\mathcal{R}$-submodule of $\mathcal{R}\otimes M.$  Since $\mathcal{R}\otimes M$ is finitely generated $\mathcal{R}$-module so $L_1(M)$ is a finitely generated $\mathcal{R}$-module.
\end{point}

\begin{point}
Superficial element\\
    \normalfont
   An element $x\in  \m$ is called $M$-superficial with respect to $\m$ if there exists $c\in \N$ such that for all $n \geq c$, $(\m^{n+1}M :_Mx)\cap \m^cM = \m^nM$. If depth $M>0$ then one can show that a $M$-superficial element is $M$-regular. Furthermore $(\m^{n+1}M :_Mx) = \m^nM$ for $n\gg0.$ Superficial elements exist if the residue field is infinite. See [\cite{JDSally}, p. 6-8] for proof of these facts when $M = A$, the same proof works in general.

    A sequence $x_1, \ldots ,x_r$  in $(A, \m)$ is said to be $M$-superficial sequence if $x_1$ is $M$-superficial and $x_i$ is $M/(x_1,\ldots ,x_{i-1})M$-superficial for $2\leq i \leq r .$
\end{point}

\begin{point}[See \cite{HCCMM}]
\label{hcs}
\normalfont
Let $(A,\m)$ be a Cohen Macaulay local ring and $M$ a finite Cohen Macaulay $A$-module of dimension $r>0.$ Let $x_1,\ldots ,x_s$ be $A\oplus M$-superficial sequence with $s\leq r$. Set $J=(x_1,\ldots ,x_s)$. The local ring $(B,\mathfrak{p})=(A/J,\m/J)$ and $B$-module $N=M/JM$ satisfy:
\begin{enumerate}
    \item $x_1,\ldots ,x_s$ is a $A\oplus M$-regular sequence.
    \item dim $M$=dim $N+s$.
    \item $N$ is a Cohen Macaulay $B$-module.
    \item $e_i(M)=e_i(N)$ for all $0\leq i\leq r-s.$

    \item $\mu(M)=\mu(N).$
    \item When $s=1$, set $x=x_1$ and $b_n(x,M)=\ell((\m^{n+1}M:x)/\m^nM)$. We have:
    \begin{enumerate}
        \item[(a)] $b_0(x,M)=0$ and $b_n(x,M)=0$ for all $n\gg0.$
        \item[(b)] $H(M,n)=\sum_{i=0}^{n}H(N,i)-b_n(x,M).$
        \item[(c)] $e_r(M)=e_r(N)-(-1)^r(\sum_{i\geq 0}b_i(x,M)).$
        \item[(d)] $x^*$ is $G(M)$-regular iff $b_n(x,M)=0$ for all $n\geq 0.$
    \end{enumerate}
    \item $e_1(M)\geq e_0(M)-\mu(M).$
    \item If $e_1(M) = e_0(M) - \mu(M)$, then $G(M)$ has minimal multiplicity and is Cohen-Macaulay.
\end{enumerate}
\end{point}

\begin{point}
Quasi-polynomial \label{qasi}\\
    \normalfont
    The function $f:\Z_{\geq 0}\to \Z_{\geq 0}$ is said to be a quasi-polynomial type with period $r>0$ if there exist polynomials $P_0,P_1,\ldots ,P_{r-1}$ such that $f(mr+i)=P_i(m)$ for all $m\gg0$ and $i=0,1,\ldots ,r-1.$
\end{point}

\begin{lemma}
\label{quaso}
    Let $f:\Z_{\geq 0}\to \Z_{\geq 0}$. Then the following are equivalent:
    \begin{enumerate}
        \item $f$ is a quasi-polynomial type with period 2.
        \item $\sum_{n\geq 0}f(n)z^n=\dfrac{h(z)}{(1-z^2)^c}$ for some $h(z)\in \Z[z]$ and $c\geq 0.$
    \end{enumerate}
    Moreover if $P_0,P_1\in \mathbb{Q}[X]$ are polynomials such that $f(2m+i)=P_i(m)$ for all $m\gg0$ and $i=0,1$ then deg $P_i\leq c-1$. Set deg $f=\text{max}\{\text{deg } P_0,\text{deg }P_1\}$
\end{lemma}

\begin{corollary}
\label{deg}
     Let $f:\Z_{\geq 0}\to \Z_{\geq 0}$. If the function $g(n)=f(n)-f(n-2)$ is a quasi-polynomial type with period 2 then so is $f$. Furthermore deg $f=\text{deg }g+1$.
\end{corollary}

\begin{point}
\label{Thm:5.2}
Eisenbud operators \\
\normalfont
Let $Q$ be a local ring and $\textbf{f}=f_1,\ldots ,f_c$ be a $Q$-regular sequence. Set $A=Q/(\textbf{f}).$ The Eisenbud operators (see \cite{Eisenbud}) are constructed as follows:

Let $\F:\cdots F_{i+2}\xrightarrow{\partial}F_{i+1}\xrightarrow{\partial}F_i\ldots$ be a complex of free $A$-modules.\\
(a) Choose a sequence of free $Q$-modules $\wt{\F_i}$ and maps $\wt{\partial}$ between them $$\wt{\F}:\cdots \wt{F}_{i+2}\xrightarrow{\wt{\partial}}\wt{F}_{i+1}\xrightarrow{\wt{\partial}}\wt{F}_i\ldots$$ such that $\F=\wt{\F}\otimes A$.\\
(b) Since $\wt{\partial}^2\equiv 0$ modulo$(f_1,\ldots ,f_c),$ we may write  $\wt{\partial}^2=\sum_{j=1}^{c} f_j\wt{t_j}$ where $\wt{t_j}:\wt{F_i}\to \wt{F}_{i-2}$ are linear maps for all $i.$\\
(c) Define, for $j=1,\ldots ,c$ the map $t_j=t_j(Q,\textbf{f},\F)$ by $t_j=\wt{t_j}\otimes A.$

The operators $t_1,\ldots ,t_c$ are called Eisenbud operators associated to $\textbf{f}.$ It can be shown that
\begin{enumerate}
    \item $t_i$ are uniquely determined up to homotopy.
    \item $t_i,t_j$ commutes up to homotopy.
\end{enumerate}
\end{point}

\begin{point}
\label{Thm:5.3}
\label{5.33}
    \normalfont
    Let $R=A[t_1,\ldots ,t_c]$ be a polynomial ring over $A$ with variable $t_1,\ldots ,t_c$ of degree $2.$ Let $M$ be an $A$-module  and let $\F$ be a free resolution of $M$. So we get well defined  maps $$t_j:\Ext_A^n(M,k)\to \Ext_A^{n+2}(M,k) \text{ for all }1\leq j\leq c \text{ and all }n.$$  This turns
    $\operatorname{Ext}_A^*(M,k)=\bigoplus_{i\geq 0}\Ext_A^i(M,k)$ into a $R$-module, where $k$ is the residue field of $A.$

    Since $\m\subseteq \text{ann}(\Ext_A^i(M,k))$ for all $i\geq 0$ we get that $\operatorname{Ext}_A^*(M,k)$ is a $S=R/\m R=k[t_1,\ldots ,t_c]$-module.
\end{point}

\begin{point}
\label{Thm:5.4}
\normalfont
    In (\cite{Gulliksen}, 3.1), Gulliksen proved that if $\text{projdim}_Q(M)$ is finite then $\operatorname{Ext}_A^*(M,N)$ is finitely generated $R$-module for all  $A$-modules $N$. In (\cite{Avramov}, 3.10), Avramov proved a converse for $N=k.$ Note that if $\text{projdim}_Q(M)$ is finite then \, $\operatorname{ Ext}_A^*(M,k)$ is finitely generated graded $S$-module. Set $\textrm{cx}_A(M)=\text{dim}_S\operatorname{Ext}_A^*(M,k)$.
\end{point}

We will repeatedly use the following result. Refer to Theorem 3.7 in \cite{cxM} for the proof.
\begin{proposition}
\label{cxM}
    Let $(A,\m,k)$ be a complete intersection ring with infinite residue field and $M$ a maximal Cohen Macaulay $A$-module of complexity $\geq 2$. Set $M_i=\text{Syz}^A_i(M)$ the $i$-th sygyzy of $M$ and $S=k[t_1,\ldots ,t_c]$ (see \ref{5.33}). Then there exists $t\in S_2$ and $n_0\in \N$ such that we have surjective maps $M_{n+2}\xrightarrow{\alpha} M_{n}$ for all $n\geq n_0,$ where $\alpha$ is an induce map of $t$ and it satisfy the following properties:
    \begin{enumerate}
        \item For a fixed $n$, the maps $\Tor^A_i(\alpha, A/\m^n)$ is surjective for all $i\geq 1.$
        \item Set $K_n(t)=\text{ker}(\alpha)$ for each $n\geq n_0.$ $\operatorname{cx}(K_n(t))=\cx(M)-1$ for all $n\geq n_0.$ So we have the following short exact sequences $0\to K_n(t)\to M_{n+2}\to M_{n}\to 0$ for all $n\gg0$ (say $n\geq c$).
    \end{enumerate}
     Further, note that $K_{n+1}(t)=\text{Syz}^A_1(K_n(t))$ for all $n\geq c.$
\end{proposition}

\section{Degree of Quasi-polynomials}
Let $F$ be a minimal resolution of $M$ over A. Set $M_i=\text{syz}_i^A(M)$ for all $i\geq 1$ and $M_0=M.$ In this section, we discuss the behavior of Hilbert coefficients of $M_i$ for all $i\gg0.$
\begin{theorem}[\cite{Avramov}, Theorem 4.1]
Let $(Q,\n)$ is regular local ring and $f_1,\ldots ,f_c \in \n^2$ be $Q$-regular sequence.  Set $(A,\m)=(Q/(f_1,\ldots ,f_c), \overline{\n})$ be a complete intersection ring. Let $M$ be a maximal Cohen Macaulay $A$-module. Then $\lim_{n\to \infty}\beta_i(M)/n^{\operatorname{cx}(M)-1}$ exists.
\end{theorem}

\begin{remark}
    If $\operatorname{cx(M)}=1$ then the function $i\mapsto \beta_i(M)$ coincides with a nonzero constant polynomial for all $i\gg0$.
\end{remark}

\begin{theorem}
\label{e_0}
    Let $(A,\m)$ be a complete intersection ring and $I$ a $\m$-primary ideal. Let $M$ be a maximal Cohen Macaulay module. Then the function $i\mapsto e^I_0(M_i)$ is a quasi-polynomial type with period $2$ and degree $\operatorname{cx}(M)-1$ for all $i\gg0.$
\end{theorem}
\begin{proof}
    We may assume $A$ is complete and the residue field of $A$ is uncountable. We will prove this by induction on $\operatorname{cx}(M).$ Assume $\operatorname{cx}(M)=1.$ We may assume $M$  has no free summand. Since $M$ is a maximal Cohen Macaulay module so $M$ has periodic resolution of period $2$, $i.e.$ $\text{Syz}^A_{2i+j}(M)=\text{Syz}^A_j(M)$ for $j=0,1$. Therefore $e^I_0(M_{2i})=e^I_0(M)\neq 0$ and $e^I_0(M_{2i+1})=e^I_0(M_1)\neq 0.$ Hence the statement is true for $\operatorname{cx}(M)=1$.

    Let us assume it is true for $\operatorname{cx}(M)=r-1.$ Let $\operatorname{cx}(M)=r.$ By \ref{cxM}, we have a short exact sequence $$0\to K_i \to M_{i+2}\to M_i\to 0$$ for all $i\gg0$ (say $i\geq j_0$) with $\operatorname{cx}(K_{j_0})=\operatorname{cx}(M)-1$. Note that $K_j\cong \text{Syz}^A_{j-j_0}(K_{j_0}).$ Since multiplicity is additive on short exact sequence we get  $e^I_0(M_{i+2})-e^I_0(M_{i})=e^I_0(K_{i})$ for all $i\gg0$. By the induction hypothesis, the function $i\mapsto e^I_0(K_{i})$ is a quasi-polynomial of degree $\cx(M)-2$ for sufficiently large $i$. Therefore, the function $i\mapsto e^I_0(M_{i})$ is a quasi-polynomial of degree $\operatorname{cx}(M)-1$ for sufficiently large $i$ (see \ref{deg}).
\end{proof}

\begin{lemma}
\label{exact}
     Let $(Q,\n)$ is regular local ring and $f_1,\ldots ,f_c \in \n^2$ be $Q$-regular sequence.  Set $(A,\m)=(Q/(f_1,\ldots ,f_c), \overline{\n})$ be a complete intersection ring with an uncountable residue field of dimension $1$. Let $M$ be a maximal Cohen Macaulay $A$-module. Then $$0\to L(K_i)\to L(M_{i+2})\to L(M_i)\to 0$$ for all $i\gg0$ (for definition of $K_n$, see \ref{cxM}).
\end{lemma}
\begin{proof}
     Let $x\in \m$ be a $M_i\oplus A$-superficial element for all $i\geq 0$. By \ref{cxM}, for each fixed $n$ we have the following surjective map
     \begin{equation}
     \label{equation}
         \operatorname{Tor}_1(M_{i+2},A/\m^n)\to \operatorname{Tor}_1(M_i,A/\m^n)\to 0
     \end{equation}
      for all $i\gg0$ (say $i\geq j_{n_0}$) and by (\cite{Quasipolynomial}, Lemma 4.1), the map  $\overline{\alpha}_x:\operatorname{Tor}_1(M_i, A/\m^n)\to  \operatorname{Tor}_1(M_i, A/\m^{n+1})$ is an isomorphism for all  $n\geq \operatorname{red}(A),$ where $\overline{\alpha}_x$ is an induced map of $\alpha_x:A/\m^n\to A/\m^{n+1}$ defined by $\alpha_x(a+\m^n)=ax+\m^{n+1}.$ For $i\geq j_0=\text{max}\{j_{n_0}, \text{ for all }1\leq n\leq \text{red}(A)\}$, we now have the following commutative diagram
\begin{equation*}
    \begin{tikzcd}
        \operatorname{Tor}_1(M_{i+2},A/\m^n) \arrow{r}{} \arrow{d}{\overline{\alpha}_x} & \operatorname{Tor}_1(M_i,A/\m^n) \arrow{r}{} \arrow{d}{\overline{\alpha}_x} & 0 \\
        \operatorname{Tor}_1(M_{i+2},A/\m^{n+1}) \arrow{r}{} & \operatorname{Tor}_1(M_i,A/\m^{n+1}) \arrow[dashed]{r}{} & 0,
    \end{tikzcd}
\end{equation*}
 since $\overline{\alpha}_x$ is isomorphic and by equation \ref{equation}, we obtained the following map $\operatorname{Tor}_1(M_{i+2},A/\m^{n+1}) \to  \operatorname{Tor}_1(M_i,A/\m^{n+1})$ is surjective. By proceeding inductively we get $$0\to L(K_i)\to L(M_{i+2})\to L(M_i)\to 0$$ for all $i\gg0.$

\end{proof}

For convenience, set deg$(0)=-\infty$.
\begin{theorem}
\label{e1}
     Let $(A,\m)$ be a complete intersection ring of codimension $c\geq 2$ and dimension $d\geq 1$. Let $M$ be a maximal Cohen Macaulay $A$-module. Set $M_i=\operatorname{syz}^A_i(M)$ for all $i\geq 0.$ Then the function $i\mapsto e_1(M_i)$ is a quasi-polynomial type with period $2$ and degree $\operatorname{cx}(M)-1$ for all $i\gg0.$
\end{theorem}
\begin{proof}
We may assume $A$ is complete and the residue field of $A$ is uncountable. Let $x_1,\ldots ,x_{d}$ be $(M_i\oplus A)$-superficial sequence for all $i\geq 0$. By  (\cite{Quasipolynomial}, Lemma 2.2,) and \ref{hcs}, we have the dimension of $M_i/(x_1,\ldots ,x_{d-1})M_i$ is one for all $i\geq 0$ and $e_1(M_i)=e_1(M_i/(x_1,\ldots ,x_{d-1}))$. So it suffices to show for $d=1.$

    Let $x\in \m\setminus \m^2$ be $(M\oplus A)$-superficial element. By \ref{hcs}, $e_1(M)\geq e_0(M)-\mu(M)\geq 0$. If $e_1(M)=0$ then $e_0(M)=\mu(M)$. This says that $M$ is a Ulrich module and $M/xM\cong k^{\mu(M)}$, where $k$ is residue field of $A.$

    Assume $\operatorname{cx}(M)=1$. Then $M$ has a periodic resolution of period $2.$ Suppose $e_1(M)=0$ then $M/xM\cong k^{\mu(M)}$. Let $F \to M$ be the minimal free resolution of $M$. Then $F/xF\to M/xM$ is a minimal free resolution of $M/xM\cong k^{\mu(M)}$, so $k$ has a periodic resolution of period $2$. Therefore by (\cite{Gulliksen}, Theorem 2.3), $A/xA$ is a hypersurface ring. As $x\in \m \setminus \m ^2$, $A$ is a hypersurface ring. This is a contradiction to the hypothesis. Similarly, we can show that $e_1(M_1)\neq 0.$ Hence the statement is true for $\operatorname{cx}(M)=1$.

     Let us assume it is true for $\operatorname{cx}(M)=r-1.$ Let $\operatorname{cx}(M)=r.$ By \ref{cxM}, we have a short exact sequence $$0\to K_i \to M_{i+2}\to M_i\to 0$$ for all $i\gg0$ (say $i\geq j_0$) with $\operatorname{cx}(K_{j_0})=\operatorname{cx}(M)-1$ and by Lemma \ref{exact}, we have the following short exact sequence $$0\to L(K_i)\to L(M_{i+2})\to L(M_i)\to 0$$ for all $i\gg0.$ This implies $e_1(M_{i+2})-e_1(M_{i})=e_1(K_{i})$ for all $i\gg0.$ By the induction hypothesis, the function $i\mapsto e_1(K_{i})$ is a quasi-polynomial of degree $\cx(M)-2$ for sufficiently large $i$. Therefore, the function $i\mapsto e_1(M_{i})$ is a quasi-polynomial of degree $\operatorname{cx}(M)-1$ for sufficiently large $i$ (see \ref{deg}).
\end{proof}

\section{Leading coefficients of quasi-polynomials}
In the previous section, we established that for a maximal Cohen-Macaulay module $M$ over ring $A$, the functions $i\mapsto e_0(M_i)$, and $i\mapsto e_1(M_i)$ are of quasi-polynomial type with period $2$ and degree $\operatorname{cx}(M)-1$ for all $i\gg0.$ Due to theorem $4.1$ in \cite{Avramov}, the function $i\mapsto \mu(M_i)$ is of quasi-polynomial type with period $2$ and degree $\operatorname{cx}(M)-1$ for all $i\gg0.$ In this section, we will explore the relationship between their leading coefficients.
\begin{lemma}

\label{cxM2}
    Let $(A,\m)$ be a complete intersection ring of codimension $c\geq 2$ and dimension $d\geq 1$. Let $M$ be a maximal Cohen Macaulay $A$-module. Then for $j=0,1$ $$\lim_{n\to \infty}\dfrac{e_1(M_{2n+j})}{n^{\operatorname{cx}(M)-1}}\geq \lim_{n\to \infty}\dfrac{e_0(M_{2n+j})}{n^{\operatorname{cx}(M)-1}}-\lim_{n\to \infty}\dfrac{\mu(M_{2n+j})}{n^{\operatorname{cx}(M)-1}},$$  where $\mu(M_n)=\text{number of minimal generators of } M_n=\beta_n(M).$ In addition, if equality holds for $\operatorname{cx}(M)=2$ then $G(K_n)$ is Cohen-Macaulay for all $n\gg0$ (for definition of $K_n$, see \ref{cxM}).
\end{lemma}
\begin{proof}
    We may assume $A$ is complete and the residue field is infinite. By (\cite{Avramov}, Theorem 4.1), Theorem \ref{e_0} and Theorem \ref{e1}, the function $i\mapsto \mu(M_i)$, $i\mapsto e_0(M_i)$ and $i\mapsto e_1(M_i)$ are quasi-polynomial type of period $2$ and degree $\operatorname{cx}(M)-1$ for all $i\gg0.$ So for $n$ large enough, let
 $$\mu(M_n)=  \left\{
\begin{array}{ll}
      \gamma n^{\operatorname{cx}(M)-1}+ \text{lower degree terms} & n \text{ even}; \\
      \gamma n^{\operatorname{cx}(M)-1}+ \text{lower degree terms} & n \text{ odd},\\
\end{array}
\right.$$
 $$e_0(M_n)=  \left\{
\begin{array}{ll}
      \alpha n^{\operatorname{cx}(M)-1}+\text{lower degree terms} & n \text{ even}; \\
      \beta n^{\operatorname{cx}(M)-1}+\text{lower degree terms} & n \text{ odd},\\
\end{array}
\right.$$
and
$$e_1(M_n)=  \left\{
\begin{array}{ll}
      an^{\operatorname{cx}(M)-1}+\text{lower degree terms} & n \text{ even}; \\
      bn^{\operatorname{cx}(M)-1}+\text{lower degree terms} & n \text{ odd}.\\
\end{array}
\right.$$
We also have $e_1(M_n)\geq e_0(M_n)-\mu(M_n)$ for all $n\geq 0$ (see \ref{hcs}). This implies
$$\lim_{n\to \infty}\dfrac{e_1(M_{2n+j})}{n^{\operatorname{cx}(M)-1}}\geq \lim_{n\to \infty}\dfrac{e_0(M_{2n+j})}{n^{\operatorname{cx}(M)-1}}-\lim_{n\to \infty}\dfrac{\mu(M_{2n+j})}{n^{\operatorname{cx}(M)-1}},$$ for $j=0,1.$ This proves the first part of the result.

Now assume $\operatorname{cx}(M)=2.$ Note that we can assume the dimension of $A$ is one (see \ref{hcs}). So by the Lemma \ref{exact}, we obtain the short exact sequence $$0\to L(K_{2n})\to L(M_{2n+2})\to L(M_{2n})\to 0$$ for all sufficiently large $n$ and this implies the following equalities.
\begin{equation*}
    \begin{split}
        \mu(M_{2n+2})-\mu(M_{2n})= & \, \mu(K_{2n}) \\  e_0(M_{2n+2})-e_0(M_{2n})= & \, e_0(K_{2n}) \\ e_1(M_{2n+2})-e_1(M_{2n})= & \,  e_1(K_{2n})
    \end{split}
\end{equation*}

Assume that for $j=0$ the equality holds. $i.e.$ $$\lim_{n\to \infty}\dfrac{e_1(M_{2n})}{n}= \lim_{n\to \infty}\dfrac{e_0(M_{2n})}{n}-\lim_{n\to \infty}\dfrac{\mu(M_{2n})}{n}.$$ Thus $a=\alpha-\gamma$ and this implies  $e_1(K_{2n}) = e_0(K_{2n}) - \mu(K_{2n})$ for all $n\gg0$. Therefore, $G(K_{2n})$ has minimal multiplicity, and hence $G(K_{2n})$ is Cohen-Macaulay (see \ref{hcs}). Similarly, by assuming the equality holds for $j = 1$, we can show that $G(K_{2n+1})$ is Cohen-Macaulay for all $n \gg 0$. This proves the result.
\end{proof}

\section{Assymptotic behaviour of $\operatorname{reg }G(M_i)$}
In this section, we aim to demonstrate that if the equality holds in Theorem \ref{cxM2} and $\operatorname{cx}(M)=2$, then the regularity of $G(M_n)$ remains bounded for all $n\gg0$ (refer to Theorem \ref{regMn}). To establish this, we require the following:
\begin{point}
    \normalfont
    Recall that $\R=\bigoplus_{n\geq 0}\m^nt^n$ is Rees ring with respect to the maximal ideal $\m$. Set $\mathfrak{M}=\m \oplus \R_+$, where $\R_+=\oplus_{n\geq 1}\m^nt^n.$ Let $E$ be a graded $\R$-module. We set $H^i(E)$ as the $i$-th graded local cohomology of $E$ with respect to $\mathfrak{M}.$ Define $a(E)=\text{sup}\{n\in \Z \,|\, H^i(E)_n\neq 0\}.$
\end{point}
\begin{point}
\label{regL}
    \normalfont
    A graded $\R$-module $E$ is considered $*$-Artinian if every descending chain of graded submodules of $E$ terminates. It can be easily proved that if $E$ is $*$-Artinian then $E_n=0$ for all $n\gg0$. Let $E$ be a finitely generated $\R$-module. Then for each $i\geq 0$, $H^i(E)$ is $*$-Artinian (see \cite{Bruns}, Proposition 3.5.4). Set $$a_i(E)=a(H^i(E)).$$ The Castelnuovo-Mumford regularity of $E$ is defined as $$\text{reg }(E)=\text{max}\{a_i(E)+i\, |\, 0\leq i\leq \text{dim }E \}.$$
\end{point}
\begin{point}
    \normalfont
    Note that we have a natural surjective homogenous homomorphism  $\phi:\R\to G(A)$ with $\phi(\mathfrak{M})=G(A)_+$, where $G(A)_+$ is the irrelevant maximal ideal of $G(A)$. So by the graded independence theorem (see \cite{SBrodman}, 13.1.6) it doesn't matter which ring we use to compute local cohomology.
\end{point}
\begin{proposition}[\cite{Part1}, 4.4]
\label{Art}
    Let $(A,\m)$ be local ring and $M$ be an $A$-module. Set $l=\text{depth }(M)$. Then  $H^i(L(M))$ is $*$-Artinian for all $0 \leq i\leq l-1$.
\end{proposition}

The following result concerns the regularity of the associated graded module of $M_n$ for sufficiently large $n$.
\begin{theorem}
\label{regMn}
     Let $(A,\m)$ be a complete intersection ring of codimension $c\geq 2$ and dimension $d\geq 1$. Let $M$ be a maximal Cohen Macaulay $A$-module. If $\operatorname{cx}(M)=2$ and $$\lim_{n\to \infty}\dfrac{e_1(M_{2n+j})}{n}= \lim_{n\to \infty}\dfrac{e_0(M_{2n+j})}{n}-\lim_{n\to \infty}\dfrac{\mu(M_{2n+j})}{n}$$ for $j=0,1$. Then $\text{reg }G(M_n)$ is bounded for all $n\gg0.$
\end{theorem}
\begin{proof}
     We may assume $A$ is complete. Due to the Proposition \ref{cxM}, we have a short exact sequence
     \begin{equation}\label{gap}
         0\to K_{n} \to M_{n+2}\to M_{n}\to 0 \text{ for all } n\gg0.
     \end{equation}

     \textbf{Claim:} $0\to L(K_{n})\to L(M_{n+2})\to L(M_{n})\to 0$ for all $n\gg0.$

Note that the statement is true for $d=1$ (see Lemma \ref{exact}). Now, assume $d\geq 2$. Equation \ref{gap} induces the following exact sequence.
 $$L_1(M_{2n})\xrightarrow{g} L(K_{2n})\to L(M_{2n+2})\to L(M_{2n})\to 0.$$ Set $U=\text{Image}(g)$. So we obtain the following exact sequence $$0\to U\to L(K_{2n})\to L(M_{2n+2})\to L(M_{2n})\to 0.$$ Since $U$ is a homomorphic image of $L_1(M_{2n})$, it is a finitely generated $\mathcal{R}$-module (see \ref{L_1M}).

Therefore, for all $j\gg0$, we have the following equality
    \begin{equation*}
        \begin{split}
            \ell(U_j)= & (e_2(K_{n})-e_2(M_{n+2})+e_2(M_{n}))\binom{j+d-2}{d-2}+\text{ lower degree terms}.
        \end{split}
    \end{equation*}
    This says that the dimension of $U$ is at most $d-1$. But by the Lemma \ref{cxM2}, $G(K_{n})$ is Cohen-Macaulay and we also have $\text{Ass}_{\mathcal{R}}(L(K_{n}))=\text{Ass}_{\mathcal{R}}(G(K_{n}))$ (see \cite{TONYT}, Proposition 5.6). Thus, $U=0$. This proves the claim.

    The short exact sequence $$0\to L(K_{n})\to L(M_{n+2})\to L(M_{n})\to 0$$ induces the following long exact sequence
      \begin{equation}
     \label{4}
         \begin{split}
             0 & \to H^0(L(K_n))\to H^0(L(M_{n+2}))\to H^0(L(M_n)\to \\
             \ldots
             \\ & \to H^{d-1}(L(K_n))\to H^{d-1}(L(M_{n+2}))\to H^{d-1}(L(M_n).
         \end{split}
     \end{equation}
    Since $G(K_n)$ is Cohen-Macaulay of dimension $r$, we have $H^i(L(K_n))=0$ for $i=0,1,\ldots ,d-1$ (see \cite{Part1}, Proposition 5.2). Therefore, for $i=0,1,\ldots ,d-2$, we have $H^i(L(M_{n+2}))\cong H^i(L(M_{n}))$ for all $n\gg0$. For $i=d-1$, we have the following descending chain
     \begin{equation*}
         H^i(L(M_{n}))\supseteq H^i(L(M_{n+2}))\supseteq \ldots
     \end{equation*}
     for all $n\gg0$. Since $H^i(L(M_n))$ is $*$-Artinian, the descending chain terminates. Therefore, there exists $n_0\in \Z$ such that $H^i(L(M_{n_0}))= H^i(L(M_{2n+n_0}))$ for all $n\geq 1$. Without loss of generality, let us assume $n_0$ is even. Then
     \begin{equation}
     \label{2}
         a_i(L(M_{2n+j}))= a_i(L(M_{n_0+j})) \text{ for all } n\geq 1, i=0,1,\ldots ,d-1 \text{ and } j=0,1.
     \end{equation}

    Note that we have the following short exact sequence: $$0\to G(M_n)\to L(M_n)\to L(M_n)(-1)\to 0$$ for all $n\geq 0.$ This induces the following long exact sequence.
     \begin{equation}
     \label{3}
         \begin{split}
             0 & \to H^0(G(M_n))\to H^0(L(M_n))\to H^0(L(M_n)(-1))\to \\
             \ldots
             \\ & \to H^{d-1}(G(M_n))\to H^{d-1}(L(M_n))\to H^{d-1}(L(M_n)(-1)).
         \end{split}
     \end{equation}

From equation \ref{3}, we obtain for all $1\leq i\leq d-1$ the following:
\begin{enumerate}
    \item[(a)] $a_0(G(M_n))\leq a_0(L(M_{n}))$,
    \item[(b)] $a_i(G(M_n))\leq \text{max}\{a_{i-1}(L(M_{n}))+1,a_i(L(M_{n})) \}$,
    \item[(c)] $a_{i-1}(L(M_{n}))+1 \leq a_i(G(M_n)).$
\end{enumerate}

Additionally, it is well known that $a_d(G(M_n))\leq \text{red}_{\m}(A)-d$ for all $n\geq 0$ (see \cite{NVT3}, 3.2). Hence, by equation \ref{2}, the regularity of $G(M_n)$ is bounded for all $n\gg0.$
\end{proof}

We recover a particular case of Theorem 4.3 and Theorem 5.3 in \cite{Quasipolynomial} as an application.

\begin{corollary}(with the same hypothesis as in \ref{regMn}) The functions $i\mapsto \text{depth }G(M_{2i})$ and \\ $i\mapsto \text{depth }G(M_{2i+1})$ are constant for $i\gg0.$
\end{corollary}
\begin{proof}
   By equation \ref{2}, we obtain that for all $n\gg0$, $a_i(L(M_{2n}))$ and $a_i(L(M_{2n+1}))$ are constants, for $i=0,1,\ldots ,r-1$. Then Proposition $5.2$ in \cite{Part1} gives the result.
\end{proof}


\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
\providecommand{\MRhref}[2]{
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}

\end{document}